Jake is johns father. The sum of their age is 44. In 14 years jake will be twice as old is john. What is both their ages
Answer by stanbon(75887)

Jake is johns father. The sum of their age is 44. In 14 years jake will be twice as old is john. What is both their ages
--------------------------
Equations:
o + a = 44
a+14 = 2(o+14)
-----
Substitute for "o" and solve for "a"::
a + 14 = 2(44-a+14)
-------
a + 14 = 2(58-a)
a + 14 = 116 - 2a
3a = 102
a = 34 years (jake's age)
----
o = 44-a = 10 years (John's age)
------------------
Cheers,
Stan H.
---------------